KjSaif (kjsaif):

The temperature started at 87°F. If the temperature went down 10°F and then up 17°F, how many degrees above or below the low is the temperature? A) -17°F B) -10°F C) 10°F D) 17°F

OpenStudy (error1603):

If your temperature goes above 87 degrees, you need to focus on bringing it down. While a prolonged 87 degree temperature will not cause lasting physical harm, if steps are not taken to reduce temperature, the fever may rise.

OpenStudy (texaschic101):

87 - 10 = 77.....77 + 17 = 94 94 - 77 = 17...answer is D....it is 17 degrees above the low (because the low is 77)

KjSaif (kjsaif):

thx

OpenStudy (texaschic101):

no problem :)